OpenStudy (anonymous):

Write each equation in Ax + By = C Form x= 2y-3 y= 2/3 x+2 2y-4 =3x

OpenStudy (anonymous):

-3 = x - 2y -2/3x + y = 2 -4 = 3x + 2y

OpenStudy (mathstudent55):

For the first one, all you need to do is subtract 2y from both sides. For the second one, subtract y from both sides. Then subtract 2 from both sides. Then switch sides. Then multiply both sides by 3. For the third one, subtract 2y from both sides and then switch sides.

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

wlcm